### v1

# BERT or FastText? A Comparative Analysis of Contextual as well as Non-Contextual Embeddings

Abstract
Natural Language Processing (NLP) for low-resource languages presents significant challenges, particularly due to the scarcity of high-quality annotated data and linguistic resources. The choice of embeddings plays a critical role in enhancing the performance of NLP tasks, such as news classification, sentiment analysis, and hate speech detection, especially for low-resource languages like Marathi. In this study, we investigate the impact of various embedding techniques—Contextual BERT-based, Non-Contextual BERT-based, and FastText-based on NLP classification tasks specific to the Marathi language. Our research includes a thorough evaluation of both compressed and uncompressed embeddings, providing a comprehensive overview of how these embeddings perform across different scenarios. Specifically, we compare two BERT model embeddings, Muril and MahaBERT, as well as two FastText model embeddings, IndicFT and MahaFT. Our evaluation includes applying embeddings to a Multiple Logistic Regression (MLR) classifier for task performance assessment, as well as TSNE visualizations to observe the spatial distribution of these embeddings. The results demonstrate that contextual embeddings outperform non-contextual embeddings. Furthermore, BERT-based non-contextual embeddings extracted from the first BERT embedding layer yield better results than FastText-based embeddings, suggesting a potential alternative to FastText embeddings.
BERT or FastText? A Comparative Analysis of Contextual as well as Non-Contextual Embeddings
Abhay Shanbhag1,3, Suramya Jadhav1,3, Amogh Thakurdesai1,3, Ridhima Sinare1,3, and Raviraj Joshi2,3 1Pune Institute of Computer Technology, Pune 2Indian Institute of Technology Madras, Chennai 3L3Cube Labs, Pune
1 Introduction
Word embedding is a way of representing words into dense vectors in a continuous space such that the vectors capture the semantic relationship between the words for the models to understand the context and meaning of the text. FastText, a context-independent method, basically captures the subword information, enabling it to learn rare words, misspelled words as well as out-of-vocabulary words. It is recognized in the NLP community for its efficient performance in tasks like text classification and sentiment analysis. Despite being relatively old, it still remains one of the most effective alternatives when performing tasks on large datasets across various languages.
BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2018), a recently emerged word embedding method, which understands the meaning of a word based on its context in a sentence. This has led to improved performance for various tasks.
FastText is widely used for low-resource languages due to its subword-based approach. On the other hand, numerous variations of BERT like IndicBERT Kakwani et al. (2020), MuRIL Khanuja et al. (2021), AfriBERT Ralethe (2020), and mBERT Devlin et al. (2018) to name a few, are available for experiments with applications involving Low Resource languages. Recent studies have experimented with both FastText and BERT for various tasks.
Experiments of D’Sa et al. (2020) demonstrated that BERT embeddings outperformed FastText for classifying English text into toxic and non-toxic. Findings of Ahmed et al. (2024) suggested that BERT embeddings outperformed those of FastText with an F1 score of 84% when evaluated for depressive post-detection in Bangla.
This paper focuses on utilizing FastText and BERT for the Marathi language for the following tasks: Sentiment Classification, 2-Class and 4-Class Hate Speech Detection, and News Article classification for headlines, long paragraphs and long documents. We construct a comprehensive analysis of FastText embeddings, IndicFT Kakwani et al. (2020) and MahaFT Joshi (2022a) embeddings and BERT embeddings, muril-base-cased Khanuja et al. (2021) and marathi-bert-v2 Joshi (2022a). Additionally, we experimented on compressed as well as contextual and non-contextual variants of the BERT-based model embeddings. This analysis revealed that contextual BERT embeddings outperformed FastText for almost all the tasks evaluated.
Section 2 provides a concise review of prior research on FastText and BERT. Section 3, includes the datasets and model embeddings that are utilized for the experiments. Section 4 presents the methodology used. Section 5 presents the results and key insights drawn from the findings along with a comparative analysis of FastText embeddings and BERT. In section 6, we conclude our discussion.
The key contributions of this work are as follows:
Conducting a comprehensive study on BERT and FastText embeddings for Marathi, a low-resource language, across diverse classification tasks, including sentiment analysis, news classification, and hate speech detection.
Evaluating the impact of embedding compression by comparing the performance of compressed and uncompressed embeddings.
Investigating the differences between contextualized and non-contextualized representations in BERT embeddings.
Type | Model | MahaSent | MahaHate | MahaNews | |||
3-class | 4-class | 2-class | SHC | LDC | LPC | ||
Contextual | MahaBERT | 82.27 | 66.8 | 85.57 | 89.83 | 93.87 | 87.78 |
MahaBERT (Compressed) | 82.89 | 66.15 | 84.37 | 89.61 | 93.53 | 87.82 | |
Muril | 81.64 | 64.55 | 84.00 | 89.54 | 93.64 | 87.33 | |
Muril (Compressed) | 81.91 | 63.2 | 83.36 | 88.38 | 93.48 | 87.45 | |
Fasttext | IndicFT | 76.4 | 58.25 | 80.13 | 85.57 | 92.15 | 79.19 |
MahaFT | 78.62 | 62.75 | 81.79 | 85.89 | 92.62 | 80.32 | |
Non-Contextual | MahaBERT | 77.56 | 66.5 | 82.64 | 86.45 | 91.69 | 81.76 |
MahaBERT (Compressed) | 76.31 | 63.9 | 81.57 | 83.85 | 91.25 | 80.08 | |
Muril | 76.58 | 65.77 | 81.76 | 85.95 | 91.61 | 81.36 | |
Muril (Compressed) | 75.16 | 63.25 | 81.44 | 82.72 | 90.39 | 79.00 |
Dataset | Subdataset | Model | Avg | Variance | Std | Test |
MahaSent | 3 Class | MahaBERT | 76.56 | 0.39843 | 0.6312 | 78.01 |
MahaBERT-Compressed | 74.42 | 0.8498 | 0.9218 | 75.51 | ||
Muril | 75.53 | 0.75268 | 0.8676 | 76.53 | ||
Muril-Compressed | 72.97 | 0.48963 | 0.6997 | 75.2 | ||
MahaFT | 77.28 | 0.38282 | 0.6187 | 78.58 | ||
MahaHate | 4 Class | MahaBERT | 64.92 | 0.25203 | 0.5020 | 66.1 |
MahaBERT-Compressed | 62.77 | 0.53875 | 0.7340 | 64.1 | ||
Muril | 63.51 | 0.35307 | 0.5942 | 65.15 | ||
Muril-Compressed | 61.22 | 0.52378 | 0.7237 | 62.9 | ||
MahaFT | 62.48 | 0.22608 | 0.4755 | 62.55 | ||
2 Class | MahaBERT | 84.23 | 0.37633 | 0.6135 | 82.53 | |
MahaBERT-Compressed | 82.3 | 0.10312 | 0.3211 | 81.41 | ||
Muril | 83.69 | 0.39397 | 0.6277 | 81.63 | ||
Muril-Compressed | 81.67 | 0.20943 | 0.4576 | 81.41 | ||
MahaFT | 83.75 | 0.52153 | 0.7222 | 82.61 | ||
MahaNews | SHC | MahaBERT | 86.66 | 0.27687 | 0.5262 | 86.64 |
MahaBERT-Compressed | 84.13 | 0.36002 | 0.6000 | 83.81 | ||
Muril | 85.7 | 0.06973 | 0.2641 | 85.66 | ||
Muril-Compressed | 82.89 | 0.11612 | 0.3408 | 82.01 | ||
MahaFT | 87.25 | 0.17873 | 0.4228 | 85.97 | ||
LDC | MahaBERT | 92.47 | 0.32565 | 0.5707 | 91.69 | |
MahaBERT-Compressed | 91.41 | 0.01637 | 0.1279 | 91.57 | ||
Muril | 92.03 | 0.19055 | 0.4365 | 91.69 | ||
Muril-Compressed | 91.04 | 0.07753 | 0.2784 | 90.39 | ||
MahaFT | 92.79 | 0.15667 | 0.3958 | 92.71 | ||
LPC | MahaBERT | 81.71 | 0.18503 | 0.4302 | 81.27 | |
MahaBERT-Compressed | 80.03 | 0.1779 | 0.4218 | 80.51 | ||
Muril | 81.19 | 0.17597 | 0.4195 | 81.4 | ||
Muril-Compressed | 78.82 | 0.14497 | 0.3807 | 79.11 | ||
MahaFT | 80.15 | 1.25257 | 1.1192 | 80.32 |
2 Literature Review
The existing literature emphasizes the superiority of BERT embeddings over traditional word embedding techniques like Word2Vec Mikolov et al. (2013) , GloVe Pennington et al. (2014), and FastText across various natural language processing (NLP) tasks. For instance, Khaled et al. (2023) compare four popular pre-trained word embeddings—Word2Vec (via Aravec Mohammad et al. (2017)), GloVe, FastText, and BERT (via ARBERTv2)—on Arabic news datasets. They highlight BERT’s superior performance, achieving over 95% accuracy due to its contextual interpretation.
Similarly, Kabullar and Türker (2022) analyzes the performance of embeddings on the AG News dataset, which includes 120K instances across four classes. They conclude that BERT outperforms other methods, achieving 90.88% accuracy. FastText, Skip-Gram, CBOW, and GloVe achieve 86.91%, 85.82%, 86.15%, and 80.86%, respectively.
While traditional embeddings perform reasonably well, the consistent dominance of BERT in complex tasks is also noted in sentiment analysis. For instance, Xie et al. (2024) explores how combining BERT and FastText embeddings enhances sentiment analysis in education, demonstrating that BERT’s contextual understanding, along with FastText’s ability to handle out-of-vocabulary words, improves generalization over unseen text.
In the domain of toxic speech classification, D’Sa et al. (2020) utilize both BERT and FastText embeddings to classify toxic comments in English, with BERT embeddings outperforming FastText. This trend continues in hate speech detection, where Rajput et al. (2021) find that neural network classifiers using BERT embeddings perform better than those with FastText embeddings alone, further supporting BERT’s effectiveness.
Additionally, Chanda (2021) assess BERT embeddings against traditional context-free methods (GloVe, Skip-Gram, and FastText) for disaster prediction, demonstrating BERT’s superior performance in combination with traditional machine learning and deep learning methods.
For low-resource languages (LRLs), Ahmed et al. (2024) examine methods like traditional TF-IDF, BERT, and FastText embeddings within a CNN-BiLSTM architecture for detecting depressive texts in Bangla. Their results show that BERT embeddings yield the highest F1 score (84%), indicating their dominance over other methods. This suggests that BERT’s efficacy extends even to LRLs.
In medical applications, Khan et al. (2024) proposes integrating BERT embeddings with SVM for prostate cancer prediction. By incorporating both numerical data and contextual information from clinical text, they achieve 95% accuracy, far outperforming the 86% accuracy achieved with numerical data alone.
Moreover, Malik et al. (2021) uses both BERT and FastText embeddings to preprocess a dataset of conversations from Twitter and Facebook. Applying various machine learning and deep learning algorithms, they find that CNN yields the best results, further demonstrating BERT’s capabilities.
Finally, while Asudani et al. (2023) offers a comprehensive analysis of traditional word embeddings alongside more advanced techniques like ELMo and BERT, providing insights into commonly used datasets and models for benchmarking, Umer et al. (2022) highlight FastText’s versatility in various domains, despite BERT’s consistently superior performance.
In conclusion, while BERT consistently outperforms other word embeddings in tasks like news classification, sentiment analysis, and hate speech detection in high-resource languages (HRLs) like English, its efficacy on LRLs remains relatively underexplored. Specifically, the comparison of BERT and FastText embeddings for NLP tasks in LRLs like news classification, sentiment analysis, and hate speech detection still warrants further investigation to assess BERT’s effectiveness across a wider range of languages and tasks.
3 Datasets and Models Used
In our research work, we used 3 Marathi datasets MahaSent: A 3-class sentiment analysis dataset Pingle et al. (2023), MahaHate: Includes both a 2-class as well as a 4-class hate classification dataset Patil et al. (2022) and MahaNews: A news categorization dataset with 12 classes Mittal et al. (2023).
We used two types of embeddings in our experiments: FastText and BERT embeddings. For FastText, we utilized both IndicFT Kakwani et al. (2020) and MahaFT Joshi (2022a) embeddings. This was because both models included a Marathi corpus as part of their training data. MahaFT, in particular, was specifically trained on a Marathi corpus, making it especially relevant for our experiments. For BERT embeddings, we primarily used two BERT-based models: MahaBERT Joshi (2022a) and MuRIL Khanuja et al. (2021). Since both models were trained on Marathi data, we selected them to compare with the FastText embeddings.
4 Methodology
For each sentence, corresponding embeddings were generated. The creation of BERT embeddings was done by first tokenizing the text using the BERT tokenizer, along with padding and truncation. The tokenized input was then passed to the model and the output of the last hidden layer of BERT was taken, which was then averaged to get contextual embeddings for every sentence. Whereas for non-contextual embeddings, the output of the first embedding layer was used.
However, for FastText, which is a non-contextual embedding by default, the process was slightly different due to the lack of a predefined vocabulary. Unlike BERT, which uses a tokenizer that can handle various Marathi words, FastText requires
the creation of a custom vocabulary. To achieve this, the training and validation datasets were concatenated and passed through a text vectorizer, which generated vectors for every word in the dataset. The vectorizer returned the vocabulary as a list of words in decreasing order of their frequency. The FastText model was then loaded using the FastText library, and for each word in the vocabulary, a word vector was retrieved to construct the embedding matrix. For each sentence, the text was split into individual words, and the corresponding embeddings were retrieved from the embedding matrix. These embeddings were then averaged to produce the final sentence embeddings.
Additionally, we experimented with compressed embeddings by reducing the dimensionality from 768 (the traditional BERT embedding dimension)
to 300. This compression was performed using
Singular Value Decomposition (SVD) to select the most relevant features, extracting the top 300 components for all the combinations of contextual as well as non-contextual for MahaBERT as well as Muril. All embeddings were then passed to a multiple logistic regression(MLR) classifier for classification into target labels. To determine the randomness, K-fold validation with 5 folds for all tasks
was performed, with the results laid out in Table 2.
4.1 Visualisation of Embeddings
5 Results
In Table 1, in the case of contextual embeddings, a decreasing trend is observed with the order: MahaBERT, Muril, MahaFT, and IndicFT. Similar results have also been observed for non-contextual embeddings. However, for non-contextual embeddings, a deviation in the above trend is observed in MahaSent and LDC datasets, where MahaFT outperforms MahaBERT.
From Table 2, we can infer that MahaSent results are inconclusive due to the high variance observed. However, we can conclude that LDC does show significant deviation from the trend, due to the relatively low variance.
Analyzing the t-SNE embedding visualizations for different datasets, BERT-based embeddings (MahaBERT and Muril) tend to form more compact clusters compared to FastText-based embeddings. Although some class overlap exists, BERT embeddings show denser class formations.
Furthermore, when evaluating the effect of compression, non-contextual embeddings tend to perform better in their uncompressed form, while no consistent trend was observed for compressed contextual embeddings.
6 Conclusion
In our research, we analyzed the effectiveness of various BERT and FastText-based embeddings on three key NLP tasks for Marathi: news classification, hate speech classification, and sentiment classification. Additionally, we examined the spatial distributions of embeddings generated for each classification task. Our study shows that MahaBERT contextual embeddings generally outperform MahaFT, Muril, and IndicNLP FastText embeddings across these tasks. This performance advantage is especially clear for contextual BERT embeddings, which show a strong trend of superiority. In contrast, non-contextual BERT embeddings generally follow this pattern but exhibit some deviations depending on the task.
7 Acknowledgement
This work was carried out under the mentorship of L3Cube, Pune. We would like to express our gratitude towards our mentor, for his continuous support and encouragement. This work is a part of the L3Cube-MahaNLP project Joshi (2022b).